# Data-driven Error Estimation: Upper Bounding Multiple Errors without Class Complexity as Input

### v3

@@ D.2 Error Estimation for Contextual Bandits
$S_{\text{def}}$ Estimate $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)=\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})$ $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{x\sim S_{err,\text{con}}}[\hat{R}_{\hat{f}}(\pi)-\hat{R}_{\hat{f}}(\hat{\pi}_{\text{con}})]$
@@
$S_{\text{err}}$ Estimate $\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{elim}}(\pi)]$ $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{con}}(\pi)]$
@@
We begin with constructing a bound $\hat{U}_{\text{elim}}$ to satisfy (27). We define our estimand analogously to the supervised learning setting described in Section 3.3, $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$. [^27] Note that while in the setting of Section Section 3.3, $\mathcal{R}$ was a measure of risk, in this setting, $R$ represents rewards. Thus, although in Section 3.3, we had $\theta_{g}=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)$, we now have $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$, i.e. the signs have switched. This is a measure of sub-optimality of policy $\hat{\pi}_{\text{elim}}$ with respect to a particular candidate policy $\pi$. However, note that while the setting in Section 3.3 was motivated by providing a bound on the estimand $\theta_{g^{*}}$, in this setting we want to bound our estimate, $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$, based on our policy evaluator $\hat{R}_{\text{elim}}(\pi)$. We define our estimate and error as
@@
To bound $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$ as needed in (27), we use the key insight that $\theta^{\text{elim}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{elim}})\geq 0$ and therefore $e^{\text{elim}}_{\pi^{*}}\geq-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$. This insight is useful for two reasons. One, it allows us to use $e^{\text{elim}}_{\pi^{*}}$ as an upper bound on our quantity of interest (and hence leverage our error estimation method). Two, this allows us to satisfy 2.3, and therefore use Corollary 2.5 to improve our bounds. Now, we use $S_{\text{err},\text{elim}}$ to estimate $\theta^{\text{elim}}(\pi)$, so we define
@@
This gives us the first bound $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq e^{\text{elim}}_{\pi^{*}}\leq\hat{\xi}_{0}^{\text{elim}}(\delta):=\max_{\pi\in\tilde{\Pi}}\hat{u}_{\text{elim}}(\pi,S_{\text{err},\text{elim}},\delta)$. Now we apply localization as in Corollary 2.5 with $c=0$, set $\Pi_{0}^{\text{elim}}=\tilde{\Pi}$, and for $k\geq 1$,
@@
2:  $\Pi_{0}^{\text{elim}}:=\{\pi\in\Pi\cup\hat{\pi}_{\text{elim}}\}$. $\hat{\xi}^{\text{elim}}_{0}(\delta):=\max_{\pi\in\Pi^{\text{elim}}_{0}}\hat{u}_{\text{elim}}(\pi,S_{\text{err}},\delta)$
@@
9:  $\hat{U}_{\text{elim}}\leftarrow\hat{\xi}_{k}^{\text{elim}}(\delta)$.
@@
We now follow a similar procedure to construct a bound $\hat{U}_{\text{con}}$ to satisfy (29). In this case, our estimand is $\theta^{\text{con}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{con}})$, which again represents the sub-optimality of policy $\hat{\pi}_{\text{con}}$ with respect to a particular candidate policy $\pi$. Our goal will be to upper bound $-\theta^{\text{con}}_{\hat{f}}(\pi^{*})$, where
@@
By Corollary 2.5, we have (39) holds, this time we set the probability to be $21-\delta/$ 2.
@@
2:  $\Pi^{\text{con}}_{0}:=\{\pi\in\Pi\cup\hat{\pi}_{\text{con}}\}$. $\hat{\xi}^{\text{con}}_{0}(\delta):=\max_{\pi\in\Pi^{\text{con}}_{0}}\hat{u}_{\text{con}}(\pi,S_{\text{err},\text{con}},\delta/3)$
@@
D.3 Contextual Arm Elimination
@@
[…]
@@
D.4 Details for Contextual Bandit Pipeline
In this Appendix we prove Theorem D.1 for the Contextual Bandit pipeline and formalize the error estimation argument in the Contextual Bandit setting.
D.4.1 Proof of Theorem 2
@@
Lemma D.3 (Conformal Uncertainty).
@@
=$\stackrel{{\scriptstyle(i)}}{{=}}\Pr_{x\sim D_{\mathcal{X}}}\Big{(}\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x))>\frac{\hat{U}_{\text{con}}}{\zeta}\Big{)}$
@@
Now, we formalize the intuition that if an arm $a$ is likely to be a good arm for context $x$ (either by belonging to both $\hat{g}(x)$ and $C(x,\beta_{\max}/\eta)$, or just to $\hat{g}(x)$), then the learner has a positive probability of pulling arm $a$ under exploration policy $p$.
Lemma D.4.
@@
$p(a|x)$$\geq\begin{cases}\frac{1-\beta_{\max}}{|C(x,\beta_{\max}/\eta)|}+\frac{\beta_{\max}}{|\hat{g}(x)|},&\text{if $a\in C(x,\beta_{\max}/\eta)$}\\ \frac{\eta}{|\hat{g}(x)|}\frac{\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)},&\text{if $a\notin C(x,\beta_{\max}/\eta)$ and $a\in\hat{g}(x)$}\\ 0,&\text{otherwise.}\end{cases}$ (48)
$\geq\begin{cases}\frac{1}{|\hat{g}(x)|},&\text{if $a\in C(x,\beta_{\max}/\eta)$}\\ \frac{\eta\hat{U}_{\text{con}}}{|\hat{g}(x)|},&\text{if $a\notin C(x,\beta_{\max}/\eta)$ and $a\in\hat{g}(x)$}\\ 0,&\text{otherwise.}\end{cases}$
@@
=$\stackrel{{\scriptstyle(iii)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I[a\in C(x,\beta/\eta)]\mbox{d}\beta$
=$\stackrel{{\scriptstyle(iv)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I\bigg{[}\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)\leq\frac{\eta\hat{U}_{\text{con}}}{\beta}\bigg{]}\mbox{d}\beta$
@@
=$\stackrel{{\scriptstyle(v)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\frac{\eta\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)}$
@@
Now, in Lemma D.5, we use the bounds derived in Lemma D.4 to establish bounds on the optimal cover $V(p,\pi^{*})$.
Lemma D.5.
@@
From Lemma D.4, we have (52) holds.
@@
Finally, we combine results from the Lemmas above to prove Theorem D.1. See D.1
@@
Since the bounds in $S$ hold with probability at least $1-\delta_{\text{prior}}$ and under this event (27) holds, we have: the bounds in $S$, $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$ hold with probability at least $31-\delta_{\text{prior}}-2\delta/$ 3. Hence, by the union bound, it is sufficient to prove (32) with probability at least $31-\delta/$ 3 under these events. Therefore moving forward, we assume $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$ and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$ hold.
Now note that $\hat{g}$ is computed in a way that satisfies (28). Since $\hat{R}_{\text{elim}}(\hat{\pi}_{\\ elim})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$, we therefore have $\hat{g}(x)\in\pi^{*}(x)$ for all $x$. Hence from $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, $\hat{g}(x)\in\pi^{*}(x)$ for all $x$, and Lemmas D.3, D.4 and D.5, we have (54) holds.
@@
Finally, by Hoeffding’s inequality, with probability at least $31-\delta/$ 3, we have
@@
D.4.2 Proof of Error Estimation for Contextual Bandits
In this section, we present the steps required to apply the Error Estimation method to the Contextual Bandit setting, and ultimately prove Corollary D.2.
@@ D.4.2 Proof of Error Estimation for Contextual Bandits
Arm Elimination Contextual Arm Sets
$S_{\text{def}}$ Estimate $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)=\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})$ $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{x\sim S_{err,\text{con}}}[\hat{R}_{\hat{f}}(\pi)-\hat{R}_{\hat{f}}(\hat{\pi}_{\text{con}})]$
$S_{\text{err}}$ Estimate $\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{elim}}(\pi)]$ $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err}}}[\theta^{\text{con}}(\pi)]$
@@
Step 1: Bounding the error in Policy Evaluation. Recall that we want to bound $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$, and since $\theta^{\text{elim}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{elim}})\geq 0$ (recall we define $\pi^{*}=\arg\max_{\pi\in\Pi}R(\pi)$, and construct $\Pi$ to include $\hat{\pi}_{\text{elim}}$), it suffices to find an upper bound for $e^{\text{elim}}_{\pi^{*}}$. Using the error estimation heuristic, we introduce $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$ as in Table 4, and write
@@
Recalling that $-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq e^{\text{elim}}_{\pi^{*}}$, this gives us the final bound
@@
Step 3: Bounding the error in the reward model. For the reward model, recall that we are interested in bounding $-\theta^{\text{con}}_{\hat{f}}(\pi^{*})$, where
@@
and use error estimation to upper bound the maximum error, and then use a single error estimate to bound the difference between this estimate and our true estimate of interest. This case is now almost identical to the arm elimination case. We similarly define $e^{\text{con}}_{\pi}=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$. Since $\theta^{\text{con}}(\pi^{*})(=R(\pi)-R(\hat{\pi}_{\text{con}}))\geq 0$, it suffices to find an upper bound for $e^{\text{con}}_{\pi^{*}}$ to bound $-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})$.
@@
    From Freedman’s inequality (Dudik et al., 2011), we have with probability at least $21-\delta/$ 2
@@
Finally, to account for having used the estimate $\hat{\theta}_{\hat{f},\text{err}_{X}}(\pi)$ rather than $\theta_{\hat{f}}(\pi)$, we bound the difference by recognizing that $\hat{f}$ is determined independently of $S_{\text{err}}$, and we can therefore use a classic Hoeffding inequality, along with a union bound argument to find that with probability at least $21-\delta/$ 2,

### v4

@@ H.2 Error Estimation for Contextual Bandits
$S_{\text{def}}$ Estimate $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi)=\hat{R}_{\text{elim}}(\pi)-\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})$ $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)=\mathbb{E}_{x\sim S_{\text{err},\text{con}}}[\hat{R}_{\hat{f}}(\pi)-\hat{R}_{\hat{f}}(\hat{\pi}_{\text{con}})]$
@@
$S_{\text{err}}$ Estimate $\hat{\theta}^{\text{elim}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err},\text{elim}}}[\theta^{\text{elim}}(\pi)]$ $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)=\mathbb{E}_{(x,a,r)\sim S_{\text{err},\text{con}}}[\theta^{\text{con}}(\pi)]$
@@
We begin with constructing a bound $\hat{U}_{\text{elim}}$ to satisfy (30). We define our estimand analogously to the supervised learning setting described in Section 3.3, $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$. [^30] Note that while in the setting of Section Section 3.3, $\mathcal{R}$ was a measure of risk, in this setting, $R$ represents rewards. Thus, although in Section 3.3, we had $\theta_{g}=\mathcal{R}(g_{\text{def}})-\mathcal{R}(g)$, we now have $\theta^{\text{elim}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{elim}})$, i.e. the signs have switched. This is a measure of sub-optimality of policy $\hat{\pi}_{\text{elim}}$ with respect to a particular candidate policy $\pi$. However, note that while the setting in Section 3.3 was motivated by providing a bound on the estimand $\theta_{g^{*}}$, in this setting we want to bound our estimate, $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$, based on our policy evaluator $\hat{R}_{\text{elim}}(\pi)$. We define our estimate and error as
@@
To bound $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$ as needed in (30), we use the key insight that $\theta^{\text{elim}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{elim}})\geq 0$ and therefore $e^{\text{elim}}_{\pi^{*}}\geq-\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$. This insight is useful for two reasons. One, it allows us to use $e^{\text{elim}}_{\pi^{*}}$ as an upper bound on our quantity of interest (and hence leverage our error estimation method). Two, this allows us to satisfy 2.3, and therefore use Corollary 2.5 to improve our bounds. Now, we use $S_{\text{err},\text{elim}}$ to estimate $\theta^{\text{elim}}(\pi)$, so we define
@@
This gives us the first bound $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq e^{\text{elim}}_{\pi^{*}}\leq\hat{\xi}_{0}^{\text{elim}}(\delta):=\max_{\pi\in\tilde{\Pi}}\hat{u}_{\text{elim}}(\pi,S_{\text{err},\text{elim}},\delta)$. Now we apply localization as in Corollary 2.5 with $c=0$, set $\Pi_{0}^{\text{elim}}=\tilde{\Pi}$, and for $k\geq 1$,
@@
[…]
@@
[…]
@@
We now follow a similar procedure to construct a bound $\hat{U}_{\text{con}}$ to satisfy (32). In this case, our estimand is $\theta^{\text{con}}(\pi)=R(\pi)-R(\hat{\pi}_{\text{con}})$, which again represents the sub-optimality of policy $\hat{\pi}_{\text{con}}$ with respect to a particular candidate policy $\pi$. Our goal will be to upper bound $\theta^{\text{con}}_{\hat{f}}(\pi^{*})$, where
@@
By Corollary 2.5, we have (42) holds, this time we set the probability to be $1-\delta/2$.
@@
[…]
@@
Hence, (30) and (32) hold.
H.3 Contextual Arm Elimination
@@
5:  Get $\text{ValidEliminator}\leftarrow\text{ValidateEliminator}(\hat{R}_{\text{elim}},\hat{\pi}_{\text{elim}},\hat{U}^{\prime},S)$.
@@
Informal guarantee: It is easy to verify that under the ideal ValidateEliminator (47), from Corollary H.2, we have (31) holds. Moreover, if the real ValidateEliminator (48) uses a holdout set $S^{\prime}$ instead of $S$, by adding a $O(\sqrt{\log(|S^{\prime}|)/n})$ term to the L.H.S. of the inequality in (48) – we can ensure (31) holds with high-probability for $(x,r)\sim D$ – for notational simplicity, we do not introduce this additional holdout set.
H.4 Details for Contextual Bandit Pipeline
In this Appendix we prove Theorem H.1 for the Contextual Bandit pipeline and formalize the error estimation argument in the Contextual Bandit setting.
H.4.1 Proof of Theorem 2
@@
Lemma H.3 (Conformal Uncertainty).
@@
$\stackrel{{\scriptstyle(i)}}{{=}}\Pr_{x\sim D_{\mathcal{X}}}\Big(\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,\pi^{*}(x))>\frac{\hat{U}_{\text{con}}}{\zeta}\Big)$
@@
Now, we formalize the intuition that if an arm $a$ is ”likely to be a good arm for context xx” (either by belonging to both $\hat{g}(x)$ and $C(x,\beta_{\max}/\eta)$, or just to $\hat{g}(x)$), then the learner has a positive probability of pulling arm $a$ under exploration policy $p$.
Lemma H.4.
@@
[…]
@@
$\stackrel{{\scriptstyle(iii)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I[a\in C(x,\beta/\eta)]\mbox{d}\beta$
$\stackrel{{\scriptstyle(iv)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\int_{0}^{1}I\bigg[\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)\leq\frac{\eta\hat{U}_{\text{con}}}{\beta}\bigg]\mbox{d}\beta$
@@
$\stackrel{{\scriptstyle(v)}}{{=}}$ $\frac{I[a\notin C(x,\beta_{\max}/\eta)]}{|\hat{g}(x)|}\frac{\eta\hat{U}_{\text{con}}}{\hat{f}(x,\hat{\pi}_{\text{con}}(x))-\hat{f}(x,a)}$
@@
Now, in Lemma H.5, we use the bounds derived in Lemma H.4 to establish bounds on the optimal cover $V(p,\pi^{*})$.
Lemma H.5.
@@
From Lemma H.4, we have (55) holds.
@@
Finally, we combine results from the Lemmas above to prove Theorem H.1. See H.1
@@
Since the bounds in $S$ hold with probability at least $1-\delta_{\text{prior}}$ and under this event (30) holds, we have: the bounds in $S$, $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$ hold with probability at least $1-\delta_{\text{prior}}-2\delta/3$. Hence, by the union bound, it is sufficient to prove (35) with probability at least $1-\delta/3$ under these events. Therefore moving forward, we assume $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$ and $\hat{R}_{\text{elim}}(\hat{\pi}_{\text{elim}})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$ hold.
Now note that $\hat{g}$ is computed in a way that satisfies (31). Since $\hat{R}_{\text{elim}}(\hat{\pi}_{\\ elim})-\hat{R}_{\text{elim}}(\pi^{*})\leq\hat{U}_{\text{elim}}$, we therefore have $\hat{g}(x)\in\pi^{*}(x)$ for all $x$. Hence from $R_{\hat{f}}(\hat{\pi}_{\text{con}})-R_{\hat{f}}(\pi^{*})\leq\hat{U}_{\text{con}}$, $\hat{g}(x)\in\pi^{*}(x)$ for all $x$, and Lemmas H.3, H.4 and H.5, we have (57) holds.
@@
Finally, by Hoeffding’s inequality, with probability at least $1-\delta/3$, we have
@@
H.4.2 Proof of Error Estimation for Contextual Bandits
In this section, we present the steps required to apply the Error Estimation method to the Contextual Bandit setting, and ultimately prove Corollary H.2.
@@ H.4.2 Proof of Error Estimation for Contextual Bandits
Step 1: Bounding the error in Policy Evaluation. Recall that we want to bound $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})$, and since $\theta^{\text{elim}}(\pi^{*})=R(\pi^{*})-R(\hat{\pi}_{\text{elim}})\geq 0$ (recall we define $\pi^{*}=\arg\max_{\pi\in\Pi}R(\pi)$, and construct $\Pi$ to include $\hat{\pi}_{\text{elim}}$), it suffices to find an upper bound for $e^{\text{elim}}_{\pi^{*}}$. Using the error estimation heuristic, we introduce $\hat{\theta}^{\text{con}}_{\text{err}}(\pi)$ as in Table 4, and write
@@
Recalling that $\hat{\theta}^{\text{elim}}_{\hat{R}}(\pi^{*})\leq e^{\text{elim}}_{\pi^{*}}$, this gives us the final bound
@@
Step 3: Bounding the error in the reward model. For the reward model, recall that we are interested in bounding $\theta^{\text{con}}_{\hat{f}}(\pi^{*})$, where
@@
and use error estimation to upper bound the maximum error, and then use a single error estimate to bound the difference between this estimate and our true estimate of interest. This case is now almost identical to the arm elimination case. We similarly define $e^{\text{con}}_{\pi}=\theta^{\text{con}}(\pi)-\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi)$. Since $\theta^{\text{con}}(\pi^{*})(=R(\pi)-R(\hat{\pi}_{\text{con}}))\geq 0$, it suffices to find an upper bound for $e^{\text{con}}_{\pi^{*}}$ to bound $\hat{\theta}^{\text{con}}_{\hat{f},\text{err}_{X}}(\pi^{*})$.
@@
    From Freedman’s inequality (Dudik et al., 2011), we have with probability at least $1-\delta/2$
@@
Finally, to account for having used the estimate $\hat{\theta}_{\hat{f},\text{err}_{X}}(\pi)$ rather than $\theta_{\hat{f}}(\pi)$, we bound the difference by recognizing that $\hat{f}$ is determined independently of $S_{\text{err}}$, and we can therefore use a classic Hoeffding inequality, along with a union bound argument to find that with probability at least $1-\delta/2$,